OpenStudy (anonymous):

find the exact area of the regions bounded by y=x^3/(sqrt(4-x^2), y=0,x=0,x=sqrt(2)

OpenStudy (anonymous):

First, the area is calculated by this integral :

OpenStudy (anonymous):

\[\int_0^{\sqrt2}\left|\frac{x^3}{\sqrt{4-x^2}}dx\right|\]